OpenStudy (anonymous):

Ken used a piece of lumber to build a bookshelf. If he made three shelves that are each 2 1/2 feet long, how long was the piece of lumber?

OpenStudy (bibby):

3 different 2.5 foot pieces can be written as 3*2.5 or 2.5+2.5+2.5

OpenStudy (anonymous):

A. 7 1/2 B. 6 1/2 C. 4 1/2 D. 5 feet

OpenStudy (bibby):

Can you not multiply 3 and 2 + 1/2?

OpenStudy (anonymous):

So you're saying 2 plus 0.5?

OpenStudy (bibby):

I shouldn't have converted to decimal \[3*2\frac{ 1 }{ 2 }=3*\frac{ 5 }{ 2 }\]

OpenStudy (anonymous):

11/2

OpenStudy (bibby):

3*5 = 15 not 11.\[\frac{ 3 }{1 }*\frac{ 5 }{2 }= \frac{ 3*5 }{1*2 } = \frac{ 15 }{2 } =7 \frac{ 1 }{2 }\]

OpenStudy (anonymous):

Okay so A

OpenStudy (anonymous):

thanks

OpenStudy (bibby):

Yeah.